OpenStudy (wesdg1978):

How do you find g '(x) from g(x)=sin x+(1/2)cot x ?

OpenStudy (anonymous):

well, the derivative of sin(x) = cos(x) and the derivative of cot(x) = -csc^2(x) So, for your function g(x) =sin (x)+(1/2)cot (x) you just separate it as 2 functions, so, g(x)=h(x)+(1/2)i(x), where h(x)=sin(x), and i(x)=cot(x) So, in the end g'(x)=h'(x)+(1/2)i'(x) and g'(x)= cos(x) - (1/2) csc^2(x)

OpenStudy (wesdg1978):

So you can just leave the 1/2 in there, you don't have to do anything with it?

OpenStudy (anonymous):

1/2 is a constant, when they're multiplying you just leave'em there, for example \[f(x)=3\cos(2x)\] then \[f'(x)=3*-sen(2x)*2=-6sen(2x)\] got it?

OpenStudy (wesdg1978):

@Umangiasd So what if g(x)= x^3cosx-13x sinx-13cosx ? How would you do that?

OpenStudy (anonymous):

First, you must apply the product rule, that says for \[f(x) = g(x) h(x)\] \[f'(x)=g'(x)h(x)+g(x)h'(x)\] in that case you must apply the rule for x^3 cos(x) and for xsinx, try and tell me what do ya get

OpenStudy (wesdg1978):

Something like this: [(3x^2)(cos x)+(-sin x)(x^3)] - [(13)(sin x) + (cos x)(13x)] - 13(-sin x)

OpenStudy (anonymous):

Excellent

OpenStudy (wesdg1978):

So is that the final answer or should I do some more simplification?

OpenStudy (anonymous):

you can put some stuff together as \[3x^2\cos(x) -x^3\sin(x) -13xcos(x)\] the 13sen(x) got cancelled

OpenStudy (wesdg1978):

Thanks for your help!